Chapter & Hw 3 0 Required information Problem 8-6A Disposal of plant assets LO C1, P1, P2 The folowing information applies to the questions displayed below Onslow Co purchases a used machine for $288.000 cash on January 2 and readies n for use the next day at a SR000 cost. On January 3, it is installed on a required operating platform costing $1,600, and it s further readed for operations The company predicts the machine will be uned for six years and have a $34.60 saoage value Deprecianon is to be changed on a straight line basis. On December 3t. at the end of its fifth year in operations, t s dsposed of Problem 8-6A Pert 2 2 Prepare journal entries to record deprecation of the machine af December a Journal entry worksheet ecord the ft yer year end adjunting eny for the depreciation expense of ateway
The following information applies to the questions displayed below Onslow Co. purchases a used machine for $288,000 cash on January 2 and readies it for use the ne cost. On January 3, it is installed on a required operating platform costing $1,600, and it is further read The company predicts the machine will be used for six years and have a $34,560 salvage value. Depr charged on a straight-line basis. On December 31, at the end of its fifth year in operations, it is dispose Problem 8-6A Part2 Print 2. Prepare journal entries to record depreciation of the machine at December 31 eferencers View transaction list Journal entry worksheet Record the year of disposal year-end adjusting entry for the depreciation expense of the used machine. Note: Enter debits before credits Debit Dec 31 Prex Type here to search Gatewa
Answer #1
Cost of machinery 297600 =288000+8000+1600
2
a
Debit Credit
Dec. 31 Depreciation expense—Machinery 43840 =(297600-34560)/6
          Accumulated depreciation—Machinery 43840
b
Dec. 31 Depreciation expense—Machinery 43840 =(297600-34560)/6
          Accumulated depreciation—Machinery 43840